# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ 5.1. Real and complex moments for the affine pure-jump stochastic volatility model
$I_{2}$ $=-\Re\Big{(}\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1\big{)}\langle\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}},\overline{u}\rangle\Big{)}\,$,
$I_{3}$ $=-\Re\Big{(}\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1\big{)}\langle\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}},\overline{u}\rangle\Big{)}.$
@@
$\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$ $\leq_{\mathcal{H}^{+}}\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\Re(u),\xi\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$ (5.8)
@@
$\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}(\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$ $=\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\int_{0}^{1}\langle\xi,u\rangle\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}\mathop{}\!\textnormal{d}s\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$
$\leq_{\mathcal{H}^{+}}\mathop{}\!\textnormal{e}^{\|\Re(u)\|_{2}}\|u\|_{\mathcal{H}^{\mathbb{C}}}\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi).$ (5.9)
@@
$I_{2}$ $\leq\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\Re(u),\xi\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|\|u\|_{\mathcal{H}^{\mathbb{C}}}+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|_{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}$
@@
where $g_{2}(u)=\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|$. Finally, for the third term in the right hand side of (5.7), it follows from inequality (5.1) and the monotonicity of $\mathcal{H}^{+}$
$I_{3}$ $\leq\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}(\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|_{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}.$
$\leq\mathop{}\!\textnormal{e}^{\|\Re(u)\|_{2}}\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi)\|\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2}.$
Hence setting $g_{3}(u)\coloneqq\mathop{}\!\textnormal{e}^{\|u\|_{\mathcal{H}^{\mathcal{C}}}}\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi)\|_{2}$, we find that (5.6) holds for the function $g=g_{1}+g_{2}+g_{3}$, which is locally Lipschitz continuous. ∎
@@
$\int_{\mathcal{H}^{+}\setminus\{0\}}\Re(e^{-\langle\xi,u\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\leq_{\mathcal{H}^{+}}\int_{\mathcal{H}^{+}\setminus\{0\}}(e^{-\langle\xi,\Re{(u)}\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}.$
@@
Note that $\mathcal{R}_{u}(t,\cdot)$ is quasi-monotone with respect to $\mathcal{H}^{+}$, for $u\in H$ (see [40, Section 6]) for the notion of quasi-monotonicity). Then by Volkmann’s comparison result [40, Satz 2] and (5.1), we have
@@
$\quad\leq\|{B}\|_{\mathcal{L}(\mathcal{H})}\|\Re(\psi_{2}(t,u))\|_{2}+\frac{1}{2}M^{2}\mathop{}\!\textnormal{e}^{2\omega t}\|D\|_{\mathcal{L}(\mathcal{H})}\|u_{1}\|_{H}$
$\qquad+M\mathop{}\!\textnormal{e}^{\omega t}\|D\|_{\mathcal{L}(\mathcal{H})}\|u_{1}\|_{H}\|\Upsilon\|_{H}$
$\qquad+\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi)\|_{2}\mathop{}\!\textnormal{e}^{\|q_{2}(t,\Re(u))\|_{2}}\|\Re{\psi_{2}(t,u)}\|_{2}$
$\qquad+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|_{2}+\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})}\mathop{}\!\textnormal{e}^{-\langle\xi,q_{2}(t,\Re(u))\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|_{2},$ (5.13)
where the constants $M\geq 1$ and $\omega\in\mathbb{R}$ are such that $\|S^{*}(t)\|_{\mathcal{L}(H)}\leq M\operatorname{e}^{\omega t}$. Hence by an application of Gronwall’s inequality we conclude that $\|\Re\psi_{2}(t,u)\|_{2}$ is bounded as long as $\|q_{2}(t,u)\|_{2}$ is bounded. Now, by Lemma 5.4, we have
@@ 5.2. Analysis of the Riccati equations in the infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, $Q$, $D\in\mathcal{S}_{1}^{+}(H)$, $\Gamma\in H$ and $\mathfrak{A}\in\mathcal{L}(H)$. Let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ and let $(\mathcal{A}^{*},\operatorname{dom}(\mathcal{A}^{*}))$ be its adjoint. Let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Then for every $u=(u_{1},u_{2})\in H\times\mathcal{S}_{n}$ there exist a positive real number $T_{\sf q_{2}}$ and a unique solution $(P_{n}(\cdot,u),q_{1}(\cdot,u),q_{2,n}(\cdot,u))$ to (2.14a)-(2.14c) on $0,T_{q_{2}})$.
@@
Standard semigroup theory ensures that for any $u_{1}\in H$ and $T>0$ the unique solution to (2.14b) is given by $q_{1,n}(t,u)=S_{t}^{*}u_{1}$ for $t\in[0,T]$. We define $\mathcal{R}_{u_{1}}^{n}(t,\cdot)\colon\mathcal{S}_{n}\to\mathcal{S}_{n}$ by
$\mathcal{R}_{u_{1}}^{n}(t,h)$ $=(\tilde{\mathcal{R}}_{n}(h)-\Pi_{n}\left(\tfrac{1}{2}\big{(}D^{1/2}S_{t}^{*}u_{1}\big{)}^{\otimes 2}-D^{1/2}S_{t}^{*}u_{1}\otimes D^{1/2}\Upsilon\right)\Pi_{n}^{*}\,$, (5.14)
@@
$\tilde{R}_{n}(h)=\Pi_{n}\left(\iota_{n}h\iota_{n}^{*}\mathfrak{A}^{*}+\mathfrak{A}\iota_{n}h\iota_{n}^{*}-\frac{1}{2}(\iota_{n}h\iota_{n}^{*}+\iota_{n}h^{T}\iota_{n}^{*})Q(\iota_{n}h\iota_{n}^{*}+\iota_{n}h^{T}\iota_{n}^{*})\right)\Pi_{n}^{*}\,.$
@@
First observe that for $u,v\in\mathcal{S}_{n}$
@@
where $u_{n}=\iota_{n}u\iota^{*}_{n}$, $v_{n}=\iota_{n}v\iota^{*}_{n}$ and $C$ is a positive constant. It follows that $\mathcal{R}_{u_{1}}(t,\cdot)$ is locally Lipschitz continuous on $\mathcal{S}_{n}$ for every $t\geq 0$ and $u_{1}\in H$. Moreover the function $\tilde{\mathcal{R}}_{n}$ is quasi-monotone with respect to $\mathcal{S}_{n}$ (see [18, Lemma 5.1]). It follows from standard ODE results (see e.g. [36, Chapter 6, Proposition 1.1]) that for every $u_{2}\in\mathcal{S}_{n}$ there exist a $T_{\sf q_{2}}>0$ and a unique solution $q_{2,n}(\cdot,u)$ of (2.14c) on $0,T_{q_{2}})$ with $T_{q_{2}}=\infty$ or
@@
By inserting $q_{2,n}(\cdot,u)$ in (2.14a) and observing that $F_{n}$ is continuous on $\mathcal{S}_{n}$, the statement follows. ∎
@@ Appendix A Laguerre polynomials and evaluation of the basis function $f_{k}$
When computing the option prices, one has to calculate the integrals w.r.t. the basis function $f_{k},k=1,\ldots,n$, see for instance, equations (4.9) and (4.17). This requires the evaluations of the basis function $f_{k}$, $k\geq 2$ which themselves are integrals of Laguerre polynomials with weight function $e^{-\frac{1}{2}(\alpha+1)}$, see equation (4.2). We start by introducing the Laguerre polynomials (see, for example, [27, Chapter V] for a good introduction to Laguerre polynomials).
@@
$L_{n}(x)=\frac{e^{x}}{n!}\frac{\mathop{}\!\textnormal{d}^{n}}{\mathop{}\!\textnormal{d}x^{n}}\left(x^{n}e^{-x}\right)\,,\qquad n\in\mathbb{N}.$
These polynomials are orthogonal with respect to the weight function $e^{-x}$ on $0,\infty)$, i.e.,
$\int_{0}^{\infty}L_{n}(x)L_{m}(x)e^{-x}\,\mathop{}\!\textnormal{d}x=\begin{cases}0&\text{if }n\neq m,\\ 1&\text{if orthonormalized},\\ \frac{1}{n!}\int_{0}^{\infty}\left[L_{n}(x)\right]^{2}e^{-x}\,\mathop{}\!\textnormal{d}x=\frac{1}{n!}&\text{(optional scaling)}.\end{cases}$
@@
$x\frac{\mathop{}\!\textnormal{d}}{\mathop{}\!\textnormal{d}x}L_{n}(x)=nL_{n}(x)-nL_{n-1}(x)\,.$ (A.2)
@@
$f_{k+3}(x)$ $=\int_{0}^{x}L_{k+1}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$=\frac{2k+1}{k+1}\int_{0}^{x}L_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s-\frac{k}{k+1}\int_{0}^{x}L_{k-1}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\qquad-\frac{1}{k+1}\int_{0}^{x}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$=\frac{2k+1}{k+1}f_{k+2}(x)-\frac{k}{k+1}f_{k+1}-\frac{1}{k+1}\int_{0}^{x}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s.$
@@
$\int_{0}^{x}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}|_{0}^{x}+\frac{2}{\alpha+1}\int_{0}^{x}[L_{k}(s)+sL_{k}^{\prime}]e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}f_{k+2}+\frac{2}{\alpha+1}\int_{0}^{x}sL_{k}^{\prime}e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}f_{k+2}$
$\qquad+\frac{2}{\alpha+1}\int_{0}^{x}[kL_{k}(s)-kL_{k-1}(s)]e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}(k+1)f_{k+2}-\frac{2}{\alpha+1}kf_{k+1}\,.$
@@
$+\frac{1}{k+1}\frac{2}{\alpha+1}\left[xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}-(k+1)f_{k+2}+kf_{k+1}\right]\,.$

### v2

@@ 5.1. Real and complex moments for the affine pure-jump stochastic volatility model
$I_{2}$ $=-\Re\Big(\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\big(\mathop{}\!\mathrm{e}^{-\langle u,\xi\rangle}-1\big)\langle\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}},\overline{u}\rangle\Big)\,$,
$I_{3}$ $=-\Re\Big(\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\big(\mathop{}\!\mathrm{e}^{-\langle u,\xi\rangle}-1\big)\langle\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}},\overline{u}\rangle\Big).$
@@
$\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\mathrm{e}^{-\langle u,\xi\rangle}\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}$ $\leq_{\mathcal{H}^{+}}\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\mathrm{e}^{-\langle\Re(u),\xi\rangle}\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}$ (5.8)
@@
$\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}(\mathop{}\!\mathrm{e}^{-\langle u,\xi\rangle}-1)\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}$ $=\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\int_{0}^{1}\langle\xi,u\rangle\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}\mathop{}\!\mathrm{d}s\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}$
$\leq_{\mathcal{H}^{+}}\mathop{}\!\mathrm{e}^{\|\Re(u)\|_{2}}\|u\|_{\mathcal{H}^{\mathbb{C}}}\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\mathrm{d}\xi).$ (5.9)
@@
$I_{2}$ $\leq\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\mathrm{e}^{-\langle\Re(u),\xi\rangle}\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}\|\|u\|_{\mathcal{H}^{\mathbb{C}}}+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|_{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}$
@@
where $g_{2}(u)=\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}\|+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|$. Finally, for the third term in the right hand side of (5.7), it follows from inequality (5.1) and the monotonicity of $\mathcal{H}^{+}$
$I_{3}$ $\leq\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}(\mathop{}\!\mathrm{e}^{-\langle u,\xi\rangle}-1)\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}\|_{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}.$
$\leq\mathop{}\!\mathrm{e}^{\|\Re(u)\|_{2}}\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\mathrm{d}\xi)\|\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2}.$
Hence setting $g_{3}(u)\coloneqq\mathop{}\!\mathrm{e}^{\|u\|_{\mathcal{H}^{\mathcal{C}}}}\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\mathrm{d}\xi)\|_{2}$, we find that (5.6) holds for the function $g=g_{1}+g_{2}+g_{3}$, which is locally Lipschitz continuous. ∎
@@
$\int_{\mathcal{H}^{+}\setminus\{0\}}\Re(\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}-1)\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}\leq_{\mathcal{H}^{+}}\int_{\mathcal{H}^{+}\setminus\{0\}}(\mathop{}\!\mathrm{e}^{-\langle\xi,\Re{(u)}\rangle}-1)\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}.$
@@
Note that $\mathcal{R}_{u}(t,\cdot)$ is quasi-monotone with respect to $\mathcal{H}^{+}$, for $u\in H$ (see [38, Section 6] for the notion of quasi-monotonicity). Then by Volkmann’s comparison result [38, Satz 2] and (5.1), we have
@@
$\quad\leq\|{B}\|_{\mathcal{L}(\mathcal{H})}\|\Re(\psi_{2}(t,u))\|_{2}+\frac{1}{2}M^{2}\mathop{}\!\mathrm{e}^{2\omega t}\|D\|_{\mathcal{L}(\mathcal{H})}\|u_{1}\|_{H}$
$\qquad+M\mathop{}\!\mathrm{e}^{\omega t}\|D\|_{\mathcal{L}(\mathcal{H})}\|u_{1}\|_{H}\|\Upsilon\|_{H}$
$\qquad+\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\mathrm{d}\xi)\|_{2}\mathop{}\!\mathrm{e}^{\|q_{2}(t,\Re(u))\|_{2}}\|\Re{\psi_{2}(t,u)}\|_{2}$
$\qquad+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|_{2}+\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})}\mathop{}\!\mathrm{e}^{-\langle\xi,q_{2}(t,\Re(u))\rangle}\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}\|_{2},$ (5.13)
where the constants $M\geq 1$ and $\omega\in\mathbb{R}$ are such that $\|S^{*}(t)\|_{\mathcal{L}(H)}\leq M\mathop{}\!\mathrm{e}^{\omega t}$. Hence by an application of Gronwall’s inequality we conclude that $\|\Re\psi_{2}(t,u)\|_{2}$ is bounded as long as $\|q_{2}(t,u)\|_{2}$ is bounded. Now, by Lemma 5.4, we have
@@ 5.2. Analysis of the Riccati equations in the infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, $Q$, $D\in\mathcal{S}_{1}^{+}(H)$, $\Gamma\in H$ and $\mathfrak{A}\in\mathcal{L}(H)$. Let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ and let $(\mathcal{A}^{*},\operatorname{dom}(\mathcal{A}^{*}))$ be its adjoint. Let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Then for every $u=(u_{1},u_{2})\in H\times\mathcal{L}_{n}$ there exist a positive real number $T_{\sf q_{2}}$ and a unique solution $(P_{n}(\cdot,u),q_{1}(\cdot,u),q_{2,n}(\cdot,u))$ to (2.14a)-(2.14c) on $0,T_{q_{2}})$.
@@
Standard semigroup theory ensures that for any $u_{1}\in H$ and $T>0$ the unique solution to (2.14b) is given by $q_{1,n}(t,u)=S_{t}^{*}u_{1}$ for $t\in[0,T]$. We define $\mathcal{R}_{u_{1}}^{n}(t,\cdot)\colon\mathcal{L}_{n}\to\mathcal{L}_{n}$ by
$\mathcal{R}_{u_{1}}^{n}(t,h)$ $=\tilde{\mathcal{R}}_{n}(h)-\Pi_{n}\left(\tfrac{1}{2}\big(D^{1/2}S_{t}^{*}u_{1}\big)^{\otimes 2}+D^{1/2}S_{t}^{*}u_{1}\otimes D^{1/2}\Upsilon\right)\Pi_{n}^{*}\,$, (5.14)
@@
$\tilde{\mathcal{R}}_{n}(h)=\Pi_{n}\left(\iota_{n}h\iota_{n}^{*}\mathfrak{A}^{*}+\mathfrak{A}\iota_{n}h\iota_{n}^{*}-\frac{1}{2}(\iota_{n}h\iota_{n}^{*}+\iota_{n}h^{T}\iota_{n}^{*})Q(\iota_{n}h\iota_{n}^{*}+\iota_{n}h^{T}\iota_{n}^{*})\right)\Pi_{n}^{*}\,.$
@@
First observe that for $u,v\in\mathcal{L}_{n}$
@@
where $u_{n}=\iota_{n}u\iota^{*}_{n}$, $v_{n}=\iota_{n}v\iota^{*}_{n}$ and $C$ is a positive constant. It follows that $\mathcal{R}_{u_{1}}^{n}(t,\cdot)$ is locally Lipschitz continuous on $\mathcal{L}_{n}$ for every $t\geq 0$ and $u_{1}\in H$. Since $\mathcal{L}_{n}\simeq\mathbb{R}^{n\times n}$ is finite-dimensional, standard ODE results (see e.g. [34, Chapter 6, Proposition 1.1]) imply that for every $u_{2}\in\mathcal{L}_{n}$ there exist a $T_{\sf q_{2}}>0$ and a unique solution $q_{2,n}(\cdot,u)$ of (2.14c) on $0,T_{q_{2}})$ with $T_{q_{2}}=\infty$ or
@@
By inserting $q_{2,n}(\cdot,u)$ in (2.14a) and observing that $F_{n}$ is continuous on $\mathcal{L}_{n}$, the statement follows. ∎
@@ Appendix A Laguerre polynomials and evaluation of the basis function $f_{k}$
When computing the option prices, one has to calculate the integrals w.r.t. the basis function $f_{k},k=1,\ldots,n$, see for instance, equations (4.9) and (4.17). This requires the evaluations of the basis function $f_{k}$, $k\geq 2$ which themselves are integrals of Laguerre polynomials with weight function $\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)}$, see equation (4.2). We start by introducing the Laguerre polynomials (see, for example, [25, Chapter V] for a good introduction to Laguerre polynomials).
@@
$L_{n}(x)=\frac{\mathop{}\!\mathrm{e}^{x}}{n!}\frac{\mathop{}\!\mathrm{d}^{n}}{\mathop{}\!\mathrm{d}x^{n}}\left(x^{n}\mathop{}\!\mathrm{e}^{-x}\right)\,,\qquad n\in\mathbb{N}.$
These polynomials are orthogonal with respect to the weight function $\mathop{}\!\mathrm{e}^{-x}$ on $0,\infty)$, i.e.,
$\int_{0}^{\infty}L_{n}(x)L_{m}(x)\mathop{}\!\mathrm{e}^{-x}\,\mathop{}\!\mathrm{d}x=\begin{cases}0&\text{if }n\neq m,\\ 1&\text{if orthonormalized},\\ \frac{1}{n!}\int_{0}^{\infty}\left[L_{n}(x)\right]^{2}\mathop{}\!\mathrm{e}^{-x}\,\mathop{}\!\mathrm{d}x=\frac{1}{n!}&\text{(optional scaling)}.\end{cases}$
@@
$x\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}x}L_{n}(x)=nL_{n}(x)-nL_{n-1}(x)\,.$ (A.2)
@@
$f_{k+3}(x)$ $=\int_{0}^{x}L_{k+1}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$=\frac{2k+1}{k+1}\int_{0}^{x}L_{k}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s-\frac{k}{k+1}\int_{0}^{x}L_{k-1}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$\qquad-\frac{1}{k+1}\int_{0}^{x}sL_{k}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$=\frac{2k+1}{k+1}f_{k+2}(x)-\frac{k}{k+1}f_{k+1}-\frac{1}{k+1}\int_{0}^{x}sL_{k}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s.$
@@
$\int_{0}^{x}sL_{k}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$\quad=-\frac{2}{\alpha+1}sL_{k}(s)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}|_{0}^{x}+\frac{2}{\alpha+1}\int_{0}^{x}[L_{k}(s)+sL_{k}^{\prime}]\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}f_{k+2}+\frac{2}{\alpha+1}\int_{0}^{x}sL_{k}^{\prime}\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}f_{k+2}$
$\qquad+\frac{2}{\alpha+1}\int_{0}^{x}[kL_{k}(s)-kL_{k-1}(s)]\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\mathrm{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}(k+1)f_{k+2}-\frac{2}{\alpha+1}kf_{k+1}\,.$
@@
$+\frac{1}{k+1}\frac{2}{\alpha+1}\left[xL_{k}(x)\mathop{}\!\mathrm{e}^{-\frac{1}{2}(\alpha+1)x}-(k+1)f_{k+2}+kf_{k+1}\right]\,.$